# HEST-1k: A Dataset for Spatial Transcriptomics and Histology Image Analysis

### v1

@@ 3.2 Histology
All tissue sections were normalized and transformed into a generic TIFF object, a pyramidal image that can easily be integrated into computational frameworks using Openslide or viewers such as QuPath Bankhead et al. (2017). In addition, we provide a contour object that delineates all the tissue regions identified in the image. We developed a robust tissue detection method where we fine-tuned a DeepLabV3 Chen et al. (2017) model with an ImageNet-pretrained ResNet50 backbone on a set of annotated segmentation regions. From the tissue segmentation, we extracted 224$\times$224-pixel patches at 20$\times$ magnification around each spot. This yielded 1.5 million valid patches for which a corresponding expression profile was derived. Such patching can readily be used for various downstream tasks, such as employed in the HEST-Benchmark or multimodal fine-tuning of foundation models for histology (Section 5 and 7).
@@ 3.3 Nuclear segmentation and classification
In addition to patching, we include nuclear segmentation that delineates each nucleus identified in all slides from HEST-1k. Specifically, we use CellViT Hörst et al. (2024), a state-of-the-art nuclear segmentation model that was trained on the PanNuke dataset Gamper et al. (2019, 2020). CellViT enables joint instance segmentation and classification of each nucleus into five classes: neoplastic epithelial, non-neoplastic epithelial, inflammatory, stromal, and necrotic. On average, we identified 55.5k nuclei per slide (standard deviation 91.3k nuclei), for a total of 59.6 million nuclei identified across all samples. Among those 17.6 million are classified as neoplastic, 21.5 million as stromal, 4.9 million as normal epithelial, 15.4 million as inflammatory, and 76 thousand as necrotic. The resulting nuclear segmentation and classification can easily be visualized using QuPath (using geojson) or loaded as Python/R objects (using json).
@@ 3.4 Gene expression
All expression data were unified in a Anndata object that can be loaded with scanpy. Anndata encodes the gene names (as var) and number of spots (as obs). Each entry represents the raw transcript counts of a gene in a given spot. No additional normalization was conducted, and we let users explore various normalization strategies based on needs, e.g., using total count normalization, log-based re-weighting, etc. In addition, we include metadata to specify the number of genes, the gene panel, and the tissue site.
To use the expression in tandem with the WSI, an alignment file describing the mapping between the image and the spots is needed. However, relying on publicly available alignment information brings three challenges: (1) most datasets report alignment with respect to a low-resolution version of the image, (2) they are not standardized, and (3) alignment quality can be low. To address these limitations, we re-aligned all samples under the same unified format between the WSI and the corresponding expression profile. To this end, we developed an automatic alignment pipeline (see Section 4) and embedded the alignment in the scanpy object.
@@ 4 HEST-Library
The HEST-Library is built around scanpy and Anndata. At its core, the HEST-Library enables (1) assembling and querying HEST-1k, and (2) running the HEST-Benchmark (see Section 5). We describe its core functionalities, in particular for unifying legacy data.
@@
Automatic alignment. Spot alignment is crucial to ensure an accurate match between the ST spots and the WSI. While software such as LoupeBrowser enables manual alignment using the fiducials (in Visium), the process remains time-consuming when processing large batches of samples. Instead, we implemented an automatic fiducial detection algorithm based on Yolov8 Redmon et al. (2016) for processing Visium samples (Appendix Figure 4). Specifically, we manually annotated 119 fiducial regions that we further augmented using tissue and fiducials mixing. We then fine-tuned Yolov8 pretrained on COCO dataset. In samples without fiducials (e.g., early versions of Spatial Transcriptomics), we realigned using the provided spot position files.
Automatic detection of image resolution. From the alignment and the spot resolution, we can infer the exact pixel resolution. We start by deriving the distance in pixel between two neighboring spots and leverage the known inter-spot distance in $\mu$m to estimate the pixel width in $\mu$m/px. For Xenium samples, we directly read the H&E alignment file which provides an affine transformation from the DAPI-stained image (with known pixel size) to the H&E image. We then compared the self-reported image resolution and our re-estimations to manually inspect and correct discrepancies.
Conversion to Anndata. ST data is provided in multiple formats, such as CSVs, MEX, TXT, h5, etc. We provide functions to unify a large set of existing formats into a AnnData object that stores the raw transcript counts as a matrix of genes by the number of spots.
Tissue segmentation and patching. We provide a tissue segmentation pipeline optimized for Visium/Xenium images. The segmented tissue can automatically be tessellated into fixed-size image patches at a predefined resolution around each spot.
Automatic HEST-1k download. To facilitate downloading part or all of the HEST-1k dataset (in total 825 GB), we implemented an easy download option where the user can specify entries of the metadata, for instance, to query all human breast cancer cases.
@@ 5 HEST-Benchmark
From HEST-1k, we curated the HEST-Benchmark, a set of ten tasks for gene expression prediction from histology in human cancer samples. The goal is two-fold: (i) benchmarking foundation models for histology under a diverse and challenging benchmark, and (ii) understanding the predictive capabilities of state-of-the-art models in predicting expression from morphology. Compared to existing tasks (e.g., Camelyon16), HEST-Benchmark brings more morphological and data acquisition diversity, and more complex tasks as expression prediction is inherently challenging.
@@ 5.1 Task definition
We define ten tasks with data from nine human cancers and ten organs (nine primary and one metastatic dataset), which include invasive ductal carcinoma (breast cancer, IDC, Task 1), prostate adenocarcinoma (prostate cancer, PRAD, Task 2), pancreatic adenocarcinoma (pancreatic cancer, PRAD, Task 3), skin cutaneous melanoma (skin cancer, SKCM, Task 4), colonic adenocarcinoma (colon cancer, COAD, Task 5), rectal adenocarcinoma (rectum cancer, READ, Task 6), clear cell renal cell carcinoma (kidney cancer, ccRCC, Task 7), hepatocellular carcinoma (liver cancer, HCC, Task 8), lung adenocarcinoma (lung cancer, LUAD, Task 9), and axillary lymph nodes in IDC (metastatic, LYM-IDC, Task 10). Additional information is provided in Appendix Table A11.
@@
5.2 Evaluating foundation models for histology on HEST-Benchmark
@@
We use the HEST-Benchmark to evaluate ten patch encoders for histology. Namely, ResNet50 (IN) Lu et al. (2020) (ImageNet pretrained), KimiaNet Riasatian et al. (2021) (pretrained with supervised subtyping classification), Ciga Ciga et al. (2022) (SimCLR pretrained on public histology data), CTransPath Wang et al. (2022) (adapted MoCov3 pretrained on TCGA and PAIP), Remedis Azizi et al. (2023) (iBOT pretrained on TCGA), Phikon Filiot et al. (2023) (iBOT pretrained on TCGA), PLIP Huang et al. (2023) (visual-language model based on CLIP pretrained on histology data from Twitter), UNI Chen et al. (2024) (DINOv2 ViT-Large pretrained on public and private histology data), CONCH Lu et al. (2024) (visual-language model using CoCa pretrained on captions from publications and educational resources), and GigaPath Xu et al. (2024) (DINOv2 ViT-Giant pretrained on proprietary data). Additional information is provided in Table A12 and Section C.3.
@@
Table 1: HEST-Benchmark results. Random Forest regression evaluation (70 trees) on publicly available state-of-the-art patch encoders. Model performance measured with Pearson correlation. We report the mean $\pm$ standard deviation computed over all folds (or patients). Best is bold, second best is underlined. Models sorted by publication date.
ResNet50 KimiaNet Ciga CTransPath Remedis Phikon PLIP UNI CONCH GigaPath
IDC 0.440 0.420 0.406 0.454 0.491 0.430 0.436 $\underline{0.502}$ $\bm{0.504}$ 0.492
±$\pm 0.03$$\pm 0.04$$\pm 0.02$$\pm 0.04$$\pm 0.06$$\pm 0.06$$\pm 0.04$$\pm 0.06$$\pm 0.05$$\pm 0.05$
PRAD 0.318 0.328 0.332 0.346 0.335 $\bm{0.377}$ 0.362 0.357 $\underline{0.373}$ 0.372
±$\pm 0.02$$\pm 0.01$$\pm 0.00$$\pm 0.01$$\pm 0.00$$\pm 0.00$$\pm 0.02$$\pm 0.00$$\pm 0.03$$\pm 0.01$
PAAD 0.389 0.410 0.397 0.406 $\bm{0.451}$ 0.372 0.392 0.424 $\underline{0.431}$ 0.425
±$\pm 0.07$$\pm 0.08$$\pm 0.07$$\pm 0.08$$\pm 0.06$$\pm 0.06$$\pm 0.09$$\pm 0.06$$\pm 0.05$$\pm 0.07$
SKCM 0.446 0.452 0.484 0.535 0.577 0.516 0.461 $\bm{0.613}$ $\underline{0.582}$ 0.541
±$\pm 0.06$$\pm 0.03$$\pm 0.01$$\pm 0.06$$\pm 0.00$$\pm 0.04$$\pm 0.06$$\pm 0.02$$\pm 0.04$$\pm 0.03$
COAD 0.107 0.080 0.102 0.123 0.125 0.137 0.112 $\bm{0.147}$ 0.124 $\underline{0.139}$
±$\pm 0.06$$\pm 0.03$$\pm 0.04$$\pm 0.05$$\pm 0.06$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.03$$\pm 0.03$
READ 0.051 0.038 0.046 0.083 0.099 0.138 0.063 $\bm{0.162}$ 0.132 $\underline{0.156}$
±$\pm 0.08$$\pm 0.05$$\pm 0.09$$\pm 0.09$$\pm 0.11$$\pm 0.10$$\pm 0.10$$\pm 0.08$$\pm 0.08$$\pm 0.08$
CCRCC 0.136 0.136 0.127 0.171 $\bm{0.200}$ 0.178 0.124 $\underline{0.186}$ 0.149 0.182
±$\pm 0.04$$\pm 0.05$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.06$$\pm 0.05$
HCC 0.034 0.028 0.045 $\bm{0.060}$ $\underline{0.059}$ 0.041 0.038 0.051 0.040 0.055
±$\pm 0.01$$\pm 0.01$$\pm 0.00$$\pm 0.00$$\pm 0.00$$\pm 0.02$$\pm 0.00$$\pm 0.00$$\pm 0.01$$\pm 0.01$
LUNG 0.497 0.507 0.515 0.531 $\bm{0.573}$ 0.541 0.533 0.511 $\underline{0.569}$ 0.547
±$\pm 0.01$$\pm 0.01$$\pm 0.02$$\pm 0.02$$\pm 0.02$$\pm 0.03$$\pm 0.02$$\pm 0.03$$\pm 0.02$$\pm 0.03$
LYMPH_IDC 0.205 0.206 0.218 0.238 0.243 0.243 0.229 0.234 $\bm{0.249}$ $\underline{0.248}$
±$\pm 0.05$$\pm 0.06$$\pm 0.07$$\pm 0.06$$\pm 0.06$$\pm 0.06$$\pm 0.06$$\pm 0.05$$\pm 0.06$$\pm 0.05$
Average 0.262 0.261 0.267 0.295 0.315 0.297 0.275 $\bm{0.319}$ 0.315 $\underline{0.316}$
@@
We learn a regression model to map model-specific patch embeddings (512 to 2048 dimensions) to the log1p-normalized expression of the top-50 highly variable genes. All tasks are evaluated using the Pearson correlation between the predicted and measured gene expression. We report mean and standard deviation across all folds (or patients). All experiments were run on a single NVIDIA 3090 GPU. Table 1 summarizes HEST-Benchmark when employing a Random Forest regression model (70 trees), and Table A13 when employing a Ridge regression model.
When using Random Forest, UNI Chen et al. (2024), GigaPath Xu et al. (2024), Remedis Azizi et al. (2023), and CONCH Lu et al. (2024) provide similar performance (0.315 to 0.319 average Pearson correlation across all tasks), outperforming all other baselines. When using a Ridge regression, CONCH Lu et al. (2024) significantly outperforms all models with an average of +5.0% absolute performance gain over the second-best model, Phikon, and 5.3% better than the third-best model, CTransPath with an average Pearson correlation of 0.316.
Supervised vs. contrastive vs. student-teacher pretraining: When employing both Random Forest or Ridge regression, supervised models (ResNet50-IN and KimiaNet) exhibit lower performance than other models. This can be attributed to (1) a lack of sufficiently large and diverse pretraining datasets and models or (2) fundamental limitations of supervised models in transferring to challenging tasks. Using Random Forest, the best student-teacher models (UNI, CONCH, GigaPath) provide similar performance to the best contrastive approach (Remedis).
Vision Transformer Base vs. Large vs. Giant: CONCH (ViT-Base, 86M parameters), UNI (ViT-Large, 307M) and GigaPath (ViT-Giant, 1.13B) provide similar performance despite GigaPath using 13$\times$ more parameters than CONCH. UNI and GigaPath have been trained in a unimodal fashion by building “views” using traditional augmentation techniques, such as color jittering and cropping. In contrast, CONCH was trained in two steps: first using unimodal iBOT pretraining, and then by fine-tuning the resulting model on image-caption pairs. Such fine-tuning step could explain its parameter-efficiency and high performance in HEST-Benchmark.
Random Forest vs. Ridge regression: The best-performing model using Random Forest (UNI, 0.319) is similar to the best-performing model using Ridge regression (CONCH, 0.316), showing the predictive capabilities of both regression models. However, Ridge regression leads to poor performance using larger models (GigaPath, Remedis, and UNI), which could be explained by the curse of dimensionality (CONCH embedding dimension is 3$\times$ smaller than GigaPath). We also emphasize that the performance could be optimized using more complex models (such as MLP) or dedicated hyper-parameter search.
Overall, HEST-Benchmark brings new insights into the performance of patch encoders for histology. We observe that (1) Random Forest regression leads to multiple models performing similarly (CONCH, UNI, GigaPath) which may suggest a form of plateauing of patch encoders, (2) CONCH is the only model that shows high performance using both Ridge and Random Forest regression models, which could be attributed to its compact size (ViT-B) and visual-language fine-tuning step, and (3) performance remains low for some tasks (e.g., READ and HCC) which implies that for certain cancer types, the morphology might not be as reflective of gene expression, as well as leaving room for follow-up works.
6 HEST for discovery
Figure 2: HEST for biomarker discovery: Analysis of an invasive ductal carcinoma Xenium sample. a. IDC Xenium samples with neoplastic nuclei overlaid in red ($n_{c}=168,033$ detected nuclei). Gray scale bar represents 2 mm. b. Heatmap of Xenium expression of gene GATA3. Blue and red values indicate above and below the mean (in white), respectively. c. Heatmap of neoplastic nuclear area. d. Four randomly selected regions with CellViT segmentation of the neoplastic nuclei. Black scale bar represents 30 $\mu$m. e., f. Correlation between area and GATA3, and minor axis length and MYBPC1.

### v2

@@ 3.2 Histology
All tissue sections were normalized and transformed into a generic TIFF object, a pyramidal image that can easily be integrated into computational frameworks using Openslide or viewers such as QuPath [bankhead2017qupath]. In addition, we provide a contour object that delineates all the tissue regions identified in the image. We developed a robust tissue vs. background detection method where we fine-tuned a DeepLabV3 [chen2017rethinking] model with an ImageNet-pretrained ResNet50 backbone on a set of annotated segmentation regions (including pen marks, fiducials, multiple stains, artifacts, etc.). From the tissue segmentation, we extracted 224$\times$224-pixel patches at 20$\times$ magnification around each spot. For Xenium samples, we generated “pseudo-Visium” spots by pooling transcripts on 55 $\times$ 55-$\mu$m patches without spacing. This yielded 2.1 million valid patches for which a corresponding expression profile was derived. Such patching can readily be used for various downstream tasks, such as employed in the HEST-Benchmark or multimodal fine-tuning of foundation models for histology (Section 5 and 7).
@@ 3.3 Nuclear segmentation and classification
In addition to patching, we include nuclear segmentation that delineates each nucleus identified in all slides from HEST-1k. We used CellViT [horst2024cellvit], a state-of-the-art nuclear segmentation model that was trained on the PanNuke dataset [gamper2019pannuke, gamper2020pannuke]. CellViT enables joint instance segmentation and classification of each nucleus into five classes: neoplastic epithelial, non-neoplastic epithelial, inflammatory, stromal, and necrotic. On average, we identified 62.1k nuclei per slide, for a total of 76.4 million nuclei identified across all samples. Among those, 17.6 million are classified as neoplastic, 21.5 million as stromal, 4.9 million as normal epithelial, 15.4 million as inflammatory, and 76 thousand as necrotic. The resulting nuclear segmentation and classification can easily be visualized using QuPath (using geojson) or loaded as Python/R objects (using JSON). For all Xenium samples, we additionally provide the nuclear and cell segmentation derived from the DAPI staining finely aligned with the H&E slide.
@@ 3.4 Gene expression
All expression data were unified in a Anndata object that can be loaded with scanpy. Anndata encodes the gene names (as var) and number of spots (as obs). Each entry represents the raw transcript counts of a gene in a given spot. No additional normalization was conducted, and we let users explore various normalization strategies based on needs, e.g., using total count normalization, log-normalization, etc. In addition, we include metadata to specify the number of genes, the gene panel, and the tissue site. For all Xenium samples, we also provide the list of all measured transcripts with their exact 2D position in the tissue (aligned with the H&E slide).
To use the expression in tandem with the WSI, an alignment file describing the mapping between the image and the spots is needed. However, relying on publicly available alignment information brings three challenges: (1) most datasets report alignment with respect to a low-resolution version of the image, (2) they are not standardized, and (3) alignment quality can be low. To address these limitations, we re-aligned all samples under the same unified format between the WSI and the corresponding expression profile. For all Visium samples, we developed an automatic alignment pipeline based on fiducial detection (see Section 4) and embedded the alignment in the scanpy object. For all Xenium samples, we used the publicly available VALIS [gatenbee2023valis] pipeline for fine-grained image registration to align the DAPI image (aligned with the transcripts by design) and the H&E slide.
@@ 4 HEST-Library
The HEST-Library is built around scanpy and Anndata. At its core, the HEST-Library enables (1) assembling and querying HEST-1k, (2) visualizing and mitigating batch effects, and (3) running the HEST-Benchmark (Section 5). We describe its core functionalities, particularly for unifying legacy data.
@@
Automatic alignment in Visium. Spot alignment is crucial to ensure an accurate match between the ST spots and the WSI. While software such as LoupeBrowser enables manual alignment using fiducials (i.e., reference markers placed at the corners of the capture area), the process remains time-consuming when processing large batches of samples. Instead, we implemented an automatic fiducial detection algorithm based on YOLOv8 [redmon2016you] for processing Visium samples (Appendix Figure 5). Specifically, we manually annotated 119 fiducial regions that we further augmented using tissue and fiducial mixing. We then fine-tuned YOLOv8 pretrained on the COCO dataset. In early versions that do not provide corner fiducials (e.g., STv1), we realigned using the provided spot position files. In Xenium, we use VALIS [gatenbee2023valis] to register the DAPI staining (aligned with the transcripts) with the H&E image.
Automatic detection of image resolution. From the alignment and the spot resolution, we can infer the exact pixel size. To this end, we compute the distance in pixel between two neighboring spots and leverage the known inter-spot distance in $\mu$m to estimate the pixel width in $\mu$m/px. For Xenium samples, we use the H&E alignment file provided as part of the assay, which provides an affine transformation from the DAPI-stained image (with known pixel size) to the H&E image. We then compared the self-reported image resolution and our re-estimations to manually inspect and correct discrepancies.
Conversion to Anndata. ST data is provided in multiple formats, such as CSV, MEX, TXT, h5, etc. We provide functions to unify a large set of existing formats into a Anndata object that stores the raw transcript counts as a matrix of genes by the number of spots, in addition to metadata about the samples (e.g., the (x,y) coordinates of each spot, the pixel resolution, etc.).
Tissue segmentation and patching. We provide a tissue segmentation pipeline optimized for Visium/Xenium images. The segmentation can then automatically tessellate the tissue into fixed-size image patches at a predefined resolution (expressed in $\mu$m/px) around each spot.
Automatic HEST-1k download. To facilitate downloading part or all of the HEST-1k dataset (over $>$1TB), we implemented an easy download option where the user can specify entries of the metadata, for instance, to query all human invasive breast cancer cases.
Batch effect visualization and mitigation. We provide functions to help visualize batch effects using dimensionality reduction techniques with user-prompted stratification (e.g., tissue site, institution, disease, etc.). In addition, we provide a wrapper of well-established batch effect mitigation strategies (namely ComBat [zhang2020combat], Harmony [korsunsky2019harmony] and matching mutual nearest neighbors [haghverdi2018mnn]), which can be applied to a list of HEST samples.
@@ 5 HEST-Benchmark
From HEST-1k, we curated the HEST-Benchmark, a set of nine tasks for gene expression prediction from histology in human cancer samples. The goal is two-fold: (i) benchmarking foundation models for histology under a diverse and challenging benchmark and (ii) understanding the predictive capabilities of state-of-the-art models in predicting expression from morphology. Compared to existing tasks (e.g., Camelyon16 [bejnordi2017diagnostic]), the HEST-Benchmark brings increased morphological diversity and more complex challenges, particularly with the inherent difficulty of expression prediction.
@@ 5.1 Task definition
We define nine tasks with data from eight human cancers and nine organs (eight primary and one metastatic dataset), which include invasive ductal carcinoma (breast cancer, IDC, Task 1), prostate adenocarcinoma (prostate cancer, PRAD, Task 2), pancreatic adenocarcinoma (pancreatic cancer, PAAD, Task 3), skin cutaneous melanoma (skin cancer, SKCM, Task 4), colonic adenocarcinoma (colon cancer, COAD, Task 5), rectal adenocarcinoma (rectum cancer, READ, Task 6), clear cell renal cell carcinoma (kidney cancer, ccRCC, Task 7), lung adenocarcinoma (lung cancer, LUAD, Task 8), and axillary lymph nodes in IDC (metastatic, LYMPH-IDC, Task 9). Additional information is provided in Appendix Table A11.
@@
5.2 Evaluating foundation model for pathology
@@
We use the HEST-Benchmark to evaluate 11 foundation models for pathology. Namely, ResNet50 (IN) [lu2020data] (ImageNet pretrained), CTransPath [wang2022transformer] (adapted MoCov3 pretrained on TCGA and PAIP), Remedis [azizi2023robust] (SimCLR [chen2020simple] pretrained on TCGA), Phikon [filiot2023scaling] (iBOT pretrained on TCGA), UNI [chen2024towards] (DINOv2 ViT-Large pretrained on internal hospital data and GTEx), CONCH [lu2024towards] (visual-language model using CoCa pretrained on captions from publications and educational resources), GigaPath [xu2024wholeslide] (DINOv2 ViT-giant pretrained on proprietary data), Virchow [vorontsov2024foundation] (DINOv2 ViT-Huge pretrained on proprietary data), Virchow 2 [Zimmermann2024Virchow2SS] (DINOv2 ViT-Huge pretrained on proprietary data), H-Optimus-0 (DINOv2 ViT-giant pretrained on proprietary data), and UNIv1.5 (DINOv2 ViT-giant pretrained on public and proprietary data). Additional information is provided in Table A12 and Section C.3.
@@
Table 1: Comparison of 11 patch encoders evaluated on the HEST-Benchmark. Reported results are based on PCA with 256 factors followed by a ridge regression. Model performance measured with Pearson correlation. Standard deviation is reported across all folds (i.e., patients). Best is bold, second best is underlined.
IDC PRAD PAAD SKCM COAD READ ccRCC LUAD LYMPH IDC Average
ResNet50 (IN) 0.4741 0.3075 0.3889 0.4822 0.2528 0.0812 0.2231 0.4917 0.2322 0.326
$\pm$ 0.047 $\pm$ 0.0309 $\pm$ 0.0754 $\pm$ 0.1141 $\pm$ 0.0372 $\pm$ 0.0517 $\pm$ 0.0554 $\pm$ 0.0119 $\pm$ 0.0491
CTransPath 0.511 0.3427 0.4378 0.5106 0.2285 0.11 0.2279 0.4985 0.2353 0.3447
$\pm$ 0.0531 $\pm$ 0.0458 $\pm$ 0.0664 $\pm$ 0.0827 $\pm$ 0.0557 $\pm$ 0.0764 $\pm$ 0.0475 $\pm$ 0.0414 $\pm$ 0.0477
Phikon 0.5327 0.342 0.4432 0.5355 0.2585 0.1517 0.2423 0.5468 0.2373 0.3656
$\pm$ 0.0914 $\pm$ 0.0767 $\pm$ 0.0684 $\pm$ 0.0549 $\pm$ 0.0056 $\pm$ 0.0822 $\pm$ 0.0263 $\pm$ 0.0045 $\pm$ 0.0457
CONCH 0.5363 0.3548 0.4475 0.5791 0.2533 0.1674 0.2179 0.5312 0.2507 0.3709
$\pm$ 0.0842 $\pm$ 0.0099 $\pm$ 0.0729 $\pm$ 0.0542 $\pm$ 0.0075 $\pm$ 0.0476 $\pm$ 0.0353 $\pm$ 0.0107 $\pm$ 0.042
REMEDIS 0.529 0.3471 0.4644 0.5818 0.2856 0.1145 0.2647 0.5336 0.2473 0.3742
$\pm$ 0.069 $\pm$ 0.0074 $\pm$ 0.0722 $\pm$ 0.0421 $\pm$ 0.02 $\pm$ 0.0987 $\pm$ 0.0539 $\pm$ 0.0326 $\pm$ 0.0585
GigaPath 0.5508 0.3708 0.4768 0.5538 0.301 0.186 0.2391 0.5399 0.2493 0.3853
$\pm$ 0.0726 $\pm$ 0.021 $\pm$ 0.0489 $\pm$ 0.0586 $\pm$ 0.0145 $\pm$ 0.0704 $\pm$ 0.0364 $\pm$ 0.0369 $\pm$ 0.0522
UNI 0.5702 0.314 0.4764 0.6254 0.263 0.1762 0.2427 0.5511 0.2565 0.3862
$\pm$ 0.0833 $\pm$ 0.0715 $\pm$ 0.0687 $\pm$ 0.0338 $\pm$ 0.0311 $\pm$ 0.0565 $\pm$ 0.0368 $\pm$ 0.0198 $\pm$ 0.0436
Virchow 0.5702 0.3309 0.4875 0.6088 0.311 0.2019 0.2637 0.5459 0.2594 0.3977
$\pm$ 0.0939 $\pm$ 0.0081 $\pm$ 0.0412 $\pm$ 0.0733 $\pm$ 0.0083 $\pm$ 0.0467 $\pm$ 0.039 $\pm$ 0.0262 $\pm$ 0.043
Virchow2 0.5922 0.3465 0.4661 0.6174 0.2578 0.2084 0.2788 0.5605 0.2582 0.3984
$\pm$ 0.0814 $\pm$ 0.029 $\pm$ 0.0676 $\pm$ 0.0174 $\pm$ 0.0189 $\pm$ 0.0502 $\pm$ 0.0516 $\pm$ 0.0172 $\pm$ 0.0296
UNIv1.5 0.5989 0.3645 0.4902 0.6401 0.2925 0.2240 0.2522 0.5586 0.2597 0.4090
$\pm$ 0.0842 $\pm$ 0.0308 $\pm$ 0.0502 $\pm$ 0.041 $\pm$ 0.0142 $\pm$ 0.0378 $\pm$ 0.04 $\pm$ 0.026 $\pm$ 0.0424
H-Optimus-0 0.5982 0.385 0.4932 0.6432 0.2991 0.2292 0.2654 0.5582 0.2595 0.4146
$\pm$ 0.0843 $\pm$ 0.0008 $\pm$ 0.0443 $\pm$ 0.0668 $\pm$ 0.0007 $\pm$ 0.041 $\pm$ 0.0309 $\pm$ 0.0324 $\pm$ 0.04
@@
We learn a regression model to map model-specific patch embeddings (512 to 2,048 dimensions) to the log1p-normalized expression of the top 50 highly variable genes. All tasks are evaluated using the Pearson correlation between the predicted and measured gene expression. We report mean and standard deviation across all folds (or patients). All experiments were run on a single NVIDIA 3090 GPU. We report performance using three downstream regression models: (i) PCA-reduced embeddings (with n=256 factors) followed by Ridge regression trained with adaptive regularization as shown in Table 1, (ii) Ridge regression model as shown in Appendix Table A13, and (iii) an XGBoost regression model with 100 estimators and a maximum depth of 3 as shown in Appendix Table A14. Our main results are reported using PCA+Ridge (i) and XGBoost (iii). Directly applying Ridge regression may unfairly penalize models with larger embedding dimensions. To guarantee a fairer and more objective comparison, we chose to utilize PCA reduction.
5.3 Scaling laws in HEST-Benchmark
Overall, H-Optimus-0 brings the best average Pearson correlation in both PCA+Ridge and XGBoost evaluation, outperforming the second-best model, UNIv1.5, by 0.56% and 0.69%, respectively. ResNet50 (IN), the only model that was not pretrained on histology images, leads to the lowest performance in both PCA+Ridge and XGBoost. Legacy domain-specific models, such as CTransPath, are outperformed by all recent models, including UNIv1.5, UNI, GigaPath, Virchow, and H-Optimus-0. The disparity between the top and bottom domain-specific models is notable, showing an absolute improvement of 7.0% for PCA+Ridge and 4.8% for XGBoost. When inspecting individual performance, we observe large differences across tasks from 0.6432 Pearson correlation in SKCM to 0.2292 in READ for H-Optimus-0 evaluated using PCA+Ridge.
Model scaling law. By inspecting the number of trainable parameters within the vision encoder for each model, we can describe how model size influences performance (measured with average Pearson correlation across all tasks, Figure 2.a). Performance increases with model size following a logarithmic scaling law (Pearson correlation of R=0.81, P-value$<$0.01). Models considered “parameter-efficient” are represented on top of the log-transformed linear regression line (e.g., CONCH, UNIv1.5, and H-Optimus-0). This observation suggests a trade-off between downstream performance and model size. Despite the observation of a model scaling law, significant variations in performance among models of identical size persist, such as between H-Optimus-0 and GigaPath, both of which are ViT-giant models with over one billion parameters.
Data scaling law. We further explored how the number of training samples used for pretraining each model (i.e., the number of image patches) affects performance. We observe that increasing the number of patches moderately correlates with the average performance (Pearson correlation of R=0.48, P-value=0.13). This correlation is weaker than model size, which we hypothesize is due to this analysis overlooking both the absolute number of WSIs used for pretraining (image patches are not independently and identically distributed per WSI) and disparities of morphological variety among WSIs (e.g., staining variation, disease diversity, artifacts, etc.).
Overall, HEST-Benchmark brings new insights into the performance of foundation models for pathology. We observe that (1) Scaling the model size strongly correlates with average performance, but the gains grow logarithmically with the number of trainable parameters. (2) Scaling the number of training patches weakly correlates with a higher performance (also on a logarithmic scale). (3) Performance remains low for some tasks (e.g., READ and ccRCC), which suggests that (i) the morphology might not be as reflective of gene expression for some cancer types or (ii) some cohorts have more noise than others (e.g., due to batch effects, low sensitivity, dropout events, or spillover between adjacent spots).
Figure 2: Scaling laws in HEST-Benchmark. a. Model scaling law comparing the number of training parameters in the vision encoder (log-scale) and the average performance on the HEST-Benchmark. Pearson correlation between parameters and performance of R=0.81 (P-value < 0.01). b. Data scaling law comparing the number of image patches used for pretraining (log-scale) and the average performance on the HEST-Benchmark. Pearson correlation between number of patches and performance of R=0.48 (P-value=0.13).
6 HEST for biomarker exploration